OpenStudy (anonymous):

let a#b#c#d be defined for all numbers by a#b#c#d=ab-cd,if x=6#3#5#4 then what is the value of 7#x#3#11?

OpenStudy (anonymous):

What happened to showing effort ;p

OpenStudy (anonymous):

First find x by the definition you gave: \[x = 6\#3\#5\#4 \implies x = 6(3) - 5(4) = 18 - 20 = -2\] Then plug in this value and do it again for those other numbers.

OpenStudy (anonymous):

then,i need to plug x=-2 into 7#x#3#7

OpenStudy (anonymous):

thanks again